A man has 3 times as much money invested in 12% bonds as he has in stocks paying 9%. If his yearly income from the investments is $2,700, what is his total investment is?
Answer by stanbon(75887)

Let amount invested at 9% be "x"
Interest on this amount is 0.09x dollars
Amount invested at 12% is "3x"
Interest on this amount is 0.12(3x)=0.36x dollars
EQUAITON:
Interest + interext = 2700 dollars
0.09x + 0.36x=2700
9x +36x=270000
45x=270000
x=6000 (amount invested at 9%)
Total investment is x+3x=4x=4(6000)=$24000.00
Cheers,
Stan H.